OpenStudy (anonymous):

Which of the following equations describes the line shown below? Check all that apply. (-2,4) (1,13)

OpenStudy (anonymous):

A. y+2=3(x-4) B. y-13=3(x-1) C. y-4=3(x-2) D.y-1=3(x-13) E. y-4=3(x+2) F. y-2=3(x-4)

OpenStudy (anonymous):

@RedoHawk what about this one

OpenStudy (anonymous):

@RedoHawk u there

OpenStudy (anonymous):

@jcpd910 can u help me

OpenStudy (anonymous):

I can help you. :)

OpenStudy (anonymous):

y+2=3(x-4) Isolate y: y = 3(x-4) - 2

OpenStudy (anonymous):

k

OpenStudy (anonymous):

(-2,4) (1,13) Those are the points you're given. Remember a point is: (x,y) So for the first equation, we can test if the x value matches the y value, like so: -2 = x y = 3(x-4) - 2 y = 3(-2 - 4) - 2 y = 3(-6) - 2 y = -18 - 2 y = -20 So when x = -2, y = -20, for this equation. So the ordered pair is: (-2, -20) which means it can't be (-2,4) So this is False. Want me to show you how to do one more?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so a will be false right

OpenStudy (anonymous):

@dan815 can u help me

OpenStudy (anonymous):

I can help.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

We do b the same way. y-13=3(x-1) y = 3(x - 1) + 13 x = -2 y = 3(-2 - 2) + 13 y = 3(-4) + 13 y = -12 + 13 y = 1 So when x = -2, y = 1 for this equation. (-2,1) that does not equal (-2, 4) so this is False.

OpenStudy (anonymous):

k what about c

OpenStudy (anonymous):

I think you should try and find that out.

OpenStudy (anonymous):

I'll help you every step of the way.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Post your work on here so I can see it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so c

OpenStudy (anonymous):

I want to see your work buddy. :)

OpenStudy (anonymous):

y-=3(x-2)

OpenStudy (anonymous):

y-4=3(x-2)

OpenStudy (anonymous):

Okay, now what?

OpenStudy (anonymous):

then y=3(x-2)-4

OpenStudy (anonymous):

Not quite, you have to add for to both sides in order to isolate y.

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

then x=1

OpenStudy (anonymous):

I want to see all of the work. :)

OpenStudy (anonymous):

y=3(x-2)-4 Now what?

OpenStudy (anonymous):

k y-4=3(x-2) then y=3(x-2)-4 then x-1

OpenStudy (anonymous):

How do you figure x = -1?

OpenStudy (anonymous):

because I subtract 3 from 2

OpenStudy (anonymous):

That's not how it works buddy. :) 3 is multiplied by the sum of x - 2.

OpenStudy (anonymous):

You can't find x with this equation. You have to sub it in with the point you're given and solve for y.

OpenStudy (anonymous):

The reason I'm using -2 for x is because the question gives you that as an x value.

OpenStudy (anonymous):

y = 3(x - 2) - 4 And then you plug -2 in for x. y = 3(-2 - 2) - 4

OpenStudy (anonymous):

ughhh this is y I hate math

OpenStudy (anonymous):

It's okay, it's difficult sometimes. :) Solve for y when x is -2 y = 3(x - 2) - 4 y = 3(-2 - 2) - 4

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

im jus thinking

OpenStudy (anonymous):

Okay, so solve this: y = 3(-2 - 2) - 4 Do you know the first step?

OpenStudy (anonymous):

i think you subtract 4 from each side

OpenStudy (anonymous):

No, you already isolated y.

OpenStudy (anonymous):

PEMDAS P - parenthesis E - exponents M - multiplication D - division A - addition S - subtraction That's the order of operations, so P comes first. Solve what's inside the parenthesis.

OpenStudy (anonymous):

y = 3(-2 - 2) - 4 P (-2 - 2)

OpenStudy (anonymous):

-2-2

OpenStudy (anonymous):

Those are minus signs. derp

OpenStudy (anonymous):

(-2-2)

OpenStudy (anonymous):

(negative two minus two)

OpenStudy (anonymous):

yea i no

OpenStudy (anonymous):

Okay, so what is it?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

Yes, so: y = 3(-2 - 2) - 4 After P y = 3(-4) - 4

OpenStudy (anonymous):

So there are no exponents, skip E, then go to M.

OpenStudy (anonymous):

u multiply 3 from -4

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

7

OpenStudy (anonymous):

No, that would be 3 + 4, not 3 * -4.

OpenStudy (anonymous):

soo sorry -12

OpenStudy (anonymous):

It's fine, so: y = -12 - 4

OpenStudy (anonymous):

Now just subtract.

OpenStudy (anonymous):

negative 12 right

OpenStudy (anonymous):

-12 - 4 is not -12. :)

OpenStudy (anonymous):

no i was saying if the 12 was negative loll

OpenStudy (anonymous):

Oh, lol yeah.

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

Yes, y = -16. So when x = -2, y = -16 for this equation. (-2, -16) Which is not: (-2, 4) So C is false. :)

OpenStudy (anonymous):

Hey, do me a favor?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

I sent you a link to my photo so you can join the domo clan. :D just put it as your picture.

OpenStudy (anonymous):

k loll

OpenStudy (anonymous):

but in this problem it has to be more than one answer

OpenStudy (anonymous):

No, it has to be 1 or more.

OpenStudy (anonymous):

Try doing D by yourself, show me each step as you complete it.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

and join the Domo clan. o.o

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@jcpd910 I here

OpenStudy (anonymous):

k Can you change your pic to the domo plz?

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

so the answers must be d,e,f

OpenStudy (anonymous):

No, we didn't check them.

OpenStudy (anonymous):

We have to check and make sure.

OpenStudy (anonymous):

k I did

OpenStudy (anonymous):

damn that domo is sexy

OpenStudy (anonymous):

mk try doing d.

OpenStudy (anonymous):

because if a,b,c are false that mean the other 3 are true

OpenStudy (anonymous):

loll

OpenStudy (anonymous):

No, there only has to be one right answer.

OpenStudy (anonymous):

Check all that apply means: Check any answer that is right Which doesn't mean three are right.

OpenStudy (anonymous):

no on this problem it says check all that apply

OpenStudy (anonymous):

I know.

OpenStudy (anonymous):

Look, watch: D.y-1=3(x-13) y = 3(x - 3) + 1 y = 3(-2 - 3) + 1 y = 3*5 + 1 y = 15 + 1 y = 16 So when x = -2 y = 16 Therefor this is false.